OpenStudy (anonymous):

What is the experimental probability that exactly 3 children in a family of 4 children will be boys? Assume that P(boy) = P(girl). (Hint: Describe your simulation and show or explain your work.)

OpenStudy (anonymous):

Number of options: 2*2*2*2=16 accepted results: B B B G B B G B B G B B G B B B = 4/16 chance = 25% chance

OpenStudy (anonymous):

thanks @Sjano

OpenStudy (anonymous):

What, no medal?